Shift report 20000704 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 04 2000 - 16:05:56 EDT

  • Next message: http account: "Shift report 20000705 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Jul  4 16:05:56 2000, Shift: 08:00-16:00
    Supervisor: RD, with: PC and many others
    Log book#2 pages 129 - 133
    
    Shift Summary:
    ------------------------------------------------------------
    When I arrived the previous shift was taking data in run 1608
    We turned on D1 and D2 at 1/6 of max field and started to take data
    in run 1609.
    
    The beam died because a "Vacuum interlock" at 9:24
    
    I had a short access as soon as RHIC went into crontolled access:
    My first task was to inspect the HV cabling of T2; I found all
    cables connected correctly.
    
    The second thing I did was to equalize the flow in MTPC1 and 2
    and when I left the IR I restored the flow to 22 cc/min in both TPCs
    
    No more beam for the rest of the shift. The operators have been trying to
    inject beam into RHIC but the machine refuses to keep it.
    
    Bjorn  pasted some plots on the log book. They show band 
    structures in the BB ZDC time correllation. These bands seem
    to appear after run 1596.
    
    D5 is on at running at 600 A (~1/4 of max field).
    
    During shift check the fan  on rack 2 NIM bin 3 was found
    unplugged. Now it blows happily.
    
    When looking at pedestals Peter found that the MultSi plot
    is identical to the one filled with data. The other plots
    for Si  look OK.
    
    C1 is now running with un updated file that has channel 4 at
    -1400 volts.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 04 2000 - 16:10:11 EDT